OpenStudy (anonymous):

Differentiate y=xcos^(-1)x-sqrt(1-x^2).

myininaya (myininaya):

This looks like an ugly one.

myininaya (myininaya):

you need to know product rule for the xcos^{-1}(x) part

myininaya (myininaya):

and chain rule for the sqrt(1-x^2) part

myininaya (myininaya):

\[\frac{d}{dx}(xcos^{-1}(x))\] so just looking at this part do you know how to differentiate this part?

OpenStudy (anonymous):

Do we need to use arccos?

myininaya (myininaya):

cos^(-1)(x) is arccos(x)

myininaya (myininaya):

\[\frac{d}{dx}\cos^{-1}(x)=\frac{-1}{\sqrt{1-x^2} }\]

myininaya (myininaya):

@tjb69812 do you know how to apply product rule to the part I mentioned above?

OpenStudy (anonymous):

f*g'(x)+f'(x)g

myininaya (myininaya):

\[\frac{d}{dx} (xcos^{-1}(x))=\cos^{-1}(x) \frac{dx}{dx}+x \frac{dcos^{-1}(x)}{dx}\]

myininaya (myininaya):

so you should be able to do this part now

myininaya (myininaya):

\[\frac{d \sqrt{1-x^2}}{dx}=\frac{d(1-x^2)^\frac{1}{2}}{dx} \]

myininaya (myininaya):

you need to know chain rule

OpenStudy (anonymous):

chain rule of the right side?

myininaya (myininaya):

\[\frac{1}{2}(1-x^2)^{\frac{1}{2}-1}(1-x^2)'\] look familiar?

OpenStudy (anonymous):

yes sir it does

zepdrix (zepdrix):

Oooo this is such a cool problem! :D Simplifies down to a really simple solution! fun stuff!

OpenStudy (anonymous):

guide me along my way!

zepdrix (zepdrix):

Oh you didn't figure this one out yet? :U

OpenStudy (anonymous):

I'm messing up somewhere it doesn't come out nicely for me :-(

zepdrix (zepdrix):

\[\Large\rm y=x \cos^{-1}x-\sqrt{1-x^2}\] \[\Large\rm \color{royalblue}{y'}=\color{royalblue}{\left(x\right)'} \cos^{-1}x+x\color{royalblue}{\left(\cos^{-1}x\right)'}-\color{royalblue}{\left(\sqrt{1-x^2}\right)'}\]Here is our setup for differentiating, product rule required. These are the things we need to differentiate, yes?

zepdrix (zepdrix):

I would recommend memorizing that the `derivative of sqrt x` is `1 over 2 sqrts of x`. It's one of those functions that shows up so often, that it's worth memorizing instead of applying power rule to.\[\Large\rm \frac{d}{dx}\sqrt{x}=\frac{1}{2\sqrt x}\]

zepdrix (zepdrix):

\[\Large\rm \color{royalblue}{y'}=\color{orangered}{\left(1\right)} \cos^{-1}x+x\color{royalblue}{\left(\cos^{-1}x\right)'}-\color{royalblue}{\left(\sqrt{1-x^2}\right)'}\]You figured out your inverse cosine derivative, yes? :O

OpenStudy (anonymous):

can you explain arccos?

zepdrix (zepdrix):

\[\Large\rm y=\cos^{-1}x\]\[\Large\rm y'=?\]We need some other approach to figure out this derivative. We'll rewrite our expression in terms of cosine,\[\Large\rm x=\cos y\]

OpenStudy (anonymous):

what rule allows that?

zepdrix (zepdrix):

You'll want to brush up on inverse functions maybe.. they're a little tricky if you haven't seen them in a while. Functions and their inverses have this property:\[\Large\rm f\left[f^{-1}(x)\right]=x\]When you take their composition, you simply get the argument back. So maybe it would make more sense to apply that to what I was trying to show you. I kind of shortcut'd a bit.\[\Large\rm y=\cos^{-1}(x)\]Apply cosine to both sides,\[\Large\rm \cos (y)=\cos\left[\cos^{-1}(x)\right]\]So our right side simplifies,\[\Large\rm \cos(y)= x\]

OpenStudy (anonymous):

yes thank you! i think i can solve from here

zepdrix (zepdrix):

Ah ok :)